Mr. Snow bought several large fries at 140 cents each. Mr. Sidney bought 2 more fries than Mr. Snow, but his where meduim fries which only cost 90 cents each. Together they spent $11. How many meduim fries did Sidney buy?
Here is what I've got so far:
x = number of fries Sidney bought
x-2 = number of fries Snow bought
140(x-2)= cost of Snow's fries
90x = cost of Sidney's fries
the correct answer is x = 6, I just can't figure out how to get it.

thanks

Answer by htmentor(1343)

Mr. Snow bought several large fries at 140 cents each. Mr. Sidney bought 2 more fries than Mr. Snow, but his where meduim fries which only cost 90 cents each. Together they spent $11. How many meduim fries did Sidney buy?
===============================================================
Let n = the number of medium fries
Then n-2 = the number of large fries
We can write the following formula for the total cost:
140(n-2) + 90n = 1100 [in cents]
Solve for n:
140n - 280 + 90n = 1100
230n = 1380
n = 6
Therefore Sidney bought 6 fries